### v1

@@ Abstract
Entity resolution (ER) is the problem of identifying and linking database records that refer to the same real-world entity. Traditional ER methods use batch processing, which becomes impractical with growing data volumes due to high computational costs and lack of real-time capabilities. In many applications, users need to resolve entities for only a small portion of their data, making full data processing unnecessary—a scenario known as “ER-on-demand”. This paper proposes FastER, an efficient ER-on-demand framework for property graphs. Our approach uses graph differential dependencies (GDDs) as a knowledge encoding language to design effective filtering mechanisms that leverage both structural and attribute semantics of graphs. We construct a blocking graph from filtered subgraphs to reduce the number of candidate entity pairs requiring comparison. Additionally, FastER incorporates Progressive Profile Scheduling (PPS), allowing the system to incrementally produce results throughout the resolution process. Extensive evaluations on multiple benchmark datasets demonstrate that FastER significantly outperforms state-of-the-art ER methods in computational efficiency and real-time processing for on-demand tasks while ensuring reliability. We make FastER publicly available at the Github link here.
@@ 1 Introduction
Entity Resolution (ER) is the problem of determining when different data records represent the same real-world entity. While critical in domains like finance, healthcare, and social networks, ER faces significant computational challenges in big data. Indeed, its quadratic time complexity means every possible entity pair must be compared in the worst case. Traditional solutions like blocking [5, 6, 1, 2] and nearest-neighbor [8, 9, 4] techniques reduce comparisons by filtering unlikely matches [3]. However, these methods often struggle with graph data, where entity relationships provide crucial context for identification. For example, in social networks and knowledge graphs, understanding connections among entities is essential for accurate duplicate detection. Moreover, conventional ER approaches may not meet the performance requirements of real-time applications where immediate results are needed. This paper studies the on-demand ER problem in property graphs, illustrated in Example 1 below.
@@
Consider a video streaming platform as a property graph: users, devices, and videos represented as nodes with multiple features; and complex relationships exist among nodes. Fig. 1 shows the ER context with six (6) key requirements. First, traditional batch ER methods are inefficient in this context as they require complete dataset processing before returning results, creating unacceptable delays for time-sensitive operations and failure to prioritize relevant entities. That is, they fail on requirements  $1,2,4$ and $5$. Second, progressive ER solutions are incapable of tackling the intricacies of property graphs (Requirement 1), and the prioritization of application-relevant subsets of the data (Requirement 4). Finally, to our knowledge, the only two existing on-demand ER solution in the literature [13, 18] are designed for relational data, thus, ill-equipped to handle Requirement $1$. We present an on-demand ER framework for property graphs, FastER, that meets all six (6) requirements. ∎
Figure 1: On-Demand ER in Property Graphs
@@
Contributions. We present a graph differential dependencies (GDDs) [34] based framework for on-demand entity resolution in property graphs. Our approach dynamically processes relevant subgraphs and progressively emits cleaned entities, enabling direct querying of dirty property graphs. More formally, the contributions of the paper are summarised as follows. 1) We leverage GDDs to encode users’ application-relevant domain knowledge, enabling relevance filtering that utilize both structural relationships and attribute semantics. 2) Further, we propose an efficient and effective blocking solution for ER-on-demand tasks that reduces computational overhead while maintaining recall. Our approach extends traditional rule-based filtering with a targeted blocking and progressive scheduling scheme, representing the first adaptation of on-demand ER to graph data. 3) We conduct extensive experiments on benchmark graphs and relational datasets. The empirical results show that our method achieves significant efficiency, offering superior real-time processing capabilities and reliability compared to existing solutions in ER-on-demand tasks.
@@
Related Work. Works related to this paper fall under three groupings below.
@@
Graph Constraints. In recent years, graph constraints have emerged as a prominent research area due to their relevance in data quality management. For instance, graph keys (GKeys) [35] leverage graph isomorphism properties to uniquely identify entities within a graph. Graph functional dependencies (GFDs) [36] apply attribute-value dependencies (similar to CFDs [37]) to the topological structures of graphs, while Graph entity dependencies (GEDs) [38] further integrate the semantics of both GFDs and GKeys. In this study, we adopt graph differential dependencies (GDDs) [34], an extension of GEDs. GDDs introduce similarity and matching semantics into the existing structure, making them suitable for defining declarative matching rules. This characteristic allows GDDs to better capture approximate matching relationships in graphs, thereby improving the accuracy and robustness of entity resolution.
@@
Progressive ER. Madhavan et al. [19] were among the first to implement a progressive data integration system in Google Base, demonstrating how web data can be integrated progressively under time and resource constraints. Since then, progressive methods have been widely applied to schema mapping [20, 21, 22, 23] and entity resolution [14, 12], especially in scenarios with limited computational capacity or time for debugging. Existing progressive ER methods employ blocking and scheduling strategies to prioritize high-probability matches for faster discovery [12, 14]. However, this approach has two key limitations. First, since result emission is not performed at the entity level, matching remains incomplete, potentially leading to incorrect intermediate results that fail to meet user requirements. Second, it lacks support for graph data structures, e.g., entity profiles generated from partial cluster matches may result in inaccurate outcomes.
@@
On-Demand ER. The most related works to this paper are in [18, 13]. Query-Driven Approach (QDA) [18] streamlines data cleaning by analyzing selection predicates within a block to reduce comparisons. However, its lack of progressive execution capability makes it unsuitable for real-time applications. On the other hand, BrewER [13] incrementally cleans data during SQL SP query evaluation using prioritized attribute comparisons. While similar to our approach, it has three key limitations: it requires computationally expensive Cartesian products to generate candidates, demands whole-dataset blocking preparation, and lacks support for graph relationships.
@@ 2.1 Basic Definitions and Notions
An entity profile for a real-world entity is a tuple $p=\langle pid,eid,type,P,R\rangle$, where $pid$ is a unique profile identifier, $eid$ is the (possibly unknown) real-world entity ID, and $type$ denotes its category (e.g., person, location). $P$ is a set of attribute-value pairs $(A,c)$, and $R$ a list of relations $(rela,pid^{\prime})$ linking the entity to other profiles. If two profiles, $p_{1},p_{2}$, represent the same entity, then $p_{1}.eid=p_{2}.eid$.
@@ 2.2 Graph Differential Dependency (GDD)
Consider $\varphi:(Q_{1}[x,x^{\prime},y],\Phi_{X}\to\Phi_{Y})$, where $\Phi_{X}=\{\delta_{Age}(x,x^{\prime})\leq 3\wedge\delta_{Ph}(x,x^{\prime})=0\}$ and $\Phi_{Y}=\{\delta_{eid}(x,x^{\prime})=0\}$. This states that if two users in $Q_{1}$ have an Age difference within $3$ and identical Phone numbers, they refer to the same real-world entity. ∎
@@ 4       $edgeWeight\leftarrow\sum_{r\in\mathcal{R}}\text{if }x,x^{\prime}\text{ satisfy}r$
6             Add edge $(x,x^{\prime})$ with $edgeWeight$ to $BlockingGraph$
@@ 12                   Compare($v,candidate$), updateMatchLists($v,candidate$)
Figure 3: The overall architecture of FastER. (a) The dirty property graph to be cleaned, (b) Subgraph obtained after graph pattern filtering, (c) Subgraph further refined by constraint-based filtering, (d)Constraits in GDD, (e) Blocking graph generated based on candidate matches and constraints, (f) The sorted profile list, (g) The neighborhood of node V11 within the same block, (h) The comparison list after processing V11 during the emission phase, (i) Comparison results of V11’s related candidates.
@@ 3.2 Constraints-Based Filtering
We define several distance constraints, which require that the differences between node pairs on certain attributes must fall within specific thresholds, such as last names, phone numbers, and ages. In this example, we define four constraints: - Lastname (LN) edit distance: $\delta_{\text{LN}}(x,x^{\prime})\leq 3$ Firstname (FN) edit distance: $\delta_{\text{FN}}(x,x^{\prime})\leq 3$ Phone (PH) edit distance: $\delta_{\text{PH}}(x,x^{\prime})\leq 3$ Age (AGE) difference: $\delta_{\text{AGE}}(x,x^{\prime})=0$. Only candidate pairs that satisfy at least one of these constraints will be retained for further processing.
@@
3.3 Block Graph Construction and PPS
Next, we construct a Blocking Graph based on the refined candidate list. In this graph, nodes represent entities, and edges represent potential matches based on their attributes. The weight of each edge is determined by the similarity between connected nodes. For simplicity and clarity, we stipulate that if two nodes satisfy one rule, the edge weight increases by one. We can define a Blocking Graph as shown in Fig. 3 e). This graph naturally divides the nodes from the original subgraph into two blocks, reducing the number of node comparisons and improving matching efficiency.
@@ 3.3 Block Graph Construction and PPS
Progressive Profile Scheduling (PPS). Once the Blocking Graph is constructed, we adopt the Progressive Profile Scheduling (PPS) technique to efficiently perform entity matching. PPS is an entity-centric scheduling method designed to prioritize nodes with the highest likelihood of matching by assessing their duplication likelihood.
@@
3.4 Time Complexity.
Given a property graph $G$ ($n$ nodes, $m$ edges) and a graph pattern $Q$ ($|V_{Q}|$ nodes, $|E_{Q}|$ edges), subgraph isomorphism complexity is $O(m\cdot|E_{Q}|)$, reducing to $O(m)$ for small patterns ($|E_{Q}|\ll m$). Using Graph Differential Dependencies (GDDs) to focus on query-relevant subgraphs reduces the candidate set size $C\ll m$, improving efficiency. For $C$ candidate node pairs, each constraint check takes $O(d)$, giving $O(C\cdot d)$. GDD constraints retain $C^{\prime}=\beta C$ ($\beta\ll 1$), reducing this to $O(\beta C)$. Blocking graph construction requires $O(C^{\prime}\cdot r)$ for edge weight computation and $O(C^{\prime}\log C^{\prime})$ for sorting. Progressive Profile Scheduling (PPS) prioritizes high-similarity pairs, reducing comparisons to $O(N\cdot k^{\prime})$ with threshold-based pruning. The overall complexity is $O(m+\beta C+C^{\prime}\log C^{\prime}+N\cdot k^{\prime}).$. In typical cases, $\beta$ and $k^{\prime}$ are small, making FastER nearly linear in $m$. A detailed derivation can be found in Appendix A.1.
@@ 4 Experimental Evaluation
Experimental Setting. All experiments were performed on a PC with a 2.40GHz Intel Xeon Silver 4210R processor and 32GB of memory, running Linux OS. The experiments were implemented in Python 3.10 and interacted with the Neo4j graph database (version 5.19.0).
Datasets. In this study, we used five graph benchmark datasets and five relational benchmark datasets (Table 1) from various domains. All datasets used in this paper are publicly available on the GitHub repository [^1] https://anonymous.4open.science/r/On_Demand_Entity_Resolution-9DFB. Since FastER is specifically designed for property graphs, we transformed the relational datasets into graph representations by leveraging their attribute relationships. The detailed conversion methods are provided in the  GitHub repository. The reported values for Nodes, Edges, Node Types, and Edge Types in the relational datasets are based on their graph-transformed statistics.
@@
Evaluation Metrics. It is important to note that in our experiments, we do not evaluate precision and F1-score. This is because FastER itself is a framework that includes blocking and scheduling components and is independent of the matching function. Therefore, the matching functions for all methods in our experiments are labeled using ground truth. Additionally, following standard practice in ER on-demand literature [13], the final matching confirmation stage guarantees a precision of 1. As a result, we exclude precision and F1-score from our evaluation. Instead, we focus on metrics such as query recall and error rate, which are discussed in detail later.
We explore the following questions in subsequent subsections:
  1. 1.
    How does FastER perform in ER-on-demand scenario (Exp-1)?
  2. 2.
    How does FastER compare to SOTA batch processing methods (Exp-2)?
  3. 3.
  4. 4.
    How does FastER perform on different components (Exp-4)?
  5. 5.
    How can users be guided to select appropriate rules (Exp-5)?

### v2

@@ Abstract
Entity resolution (ER) is the problem of identifying and linking database records that refer to the same real-world entity. Traditional ER methods use batch processing, which becomes impractical with growing data volumes due to high computational costs and a lack of real-time capabilities. In many applications, users need to resolve entities for only a small portion of their data, making full data processing unnecessary—a scenario known as “ER-on-demand”. This paper proposes FastER, an efficient ER-on-demand framework for property graphs. Our approach uses graph differential dependencies (GDD) as a knowledge-encoding language to design effective filtering mechanisms that leverage both the structural and attribute semantics of graphs. We construct a blocking graph from the filtered subgraphs to reduce the number of candidate entity pairs that require comparison. Additionally, FastER incorporates Progressive Profile Scheduling (PPS), allowing the system to incrementally produce results throughout the resolution process. Extensive evaluations on multiple benchmark datasets demonstrate that FastER significantly outperforms state-of-the-art ER methods in computational efficiency and real-time processing for on-demand tasks, without compromising quality or reliability. We make FastER publicly available at the Github link here.
@@ 1 Introduction
Entity Resolution (ER) is the problem of determining when different data records represent the same real-world entity. While critical in domains like finance, healthcare, and social networks, ER faces significant computational challenges in
Figure 1: On-Demand ER in Property Graphs
big data. Indeed, its quadratic time complexity means every possible entity pair must be compared in the worst case. Traditional solutions like blocking [5, 6, 1, 2] and nearest-neighbor [7, 8, 4] techniques reduce comparisons by filtering unlikely matches [3]. However, these methods often struggle with graph data, where entity relationships provide crucial context for identification. For example, in social networks and knowledge graphs, understanding connections among entities is essential for accurate duplicate detection. Moreover, conventional ER approaches may not meet the performance requirements of real-time applications where immediate results are needed. This paper studies the on-demand ER problem in property graphs, illustrated in Example 1 below.
@@
Consider a video streaming platform modeled as a property graph, where users, devices, and videos are represented as nodes with rich attributes, and complex relationships exist among them. Fig. 1 outlines six key requirements for entity resolution (ER) in this scenario. First, batch-based ER methods are inefficient, especially when dealing with large-scale data, as they require processing the entire dataset before returning results. This makes it difficult to meet the responsiveness needed in time-sensitive applications and to prioritize relevant entities (Requirements 1, 2, and 4). Matching functions, such as neural or knowledge-based models, are often computationally expensive (Requirement 6), making pre-filtering essential for reducing unnecessary invocations. Moreover, interpretability is a crucial requirement in many real-world ER applications (Requirement 5). Second, progressive ER methods struggle with the structural complexity of property graphs (Requirement 1) and are unable to focus on application-specific subgraphs (Requirement 4). Existing on-demand ER solutions [10, 14, 17] are primarily designed for relational data and are not well-suited to the structural needs of graph-based scenarios (Requirement 1). We propose FastER, an on-demand ER framework for property graphs that satisfies all six requirements outlined in Fig. 1. ∎
1.0.1 Contributions.
@@
We present a graph differential dependencies (GDDs) [21] based framework for on-demand entity resolution in property graphs. Our approach dynamically processes relevant subgraphs and progressively emits cleaned entities, enabling direct querying of dirty property graphs. More formally, the contributions of the paper are summarised as follows: 1) We leverage GDDs to encode users’ application-relevant domain knowledge, enabling relevance filtering that utilizes both structural relationships and attribute semantics. 2) Further, we propose an efficient and effective blocking solution for ER-on-demand tasks that reduces computational overhead while maintaining recall. Our approach extends traditional rule-based filtering with a targeted blocking scheme and a progressive scheduling scheme, representing the first adaptation of on-demand ER to graph data. 3) We conduct extensive experiments on benchmark graphs and relational datasets. The empirical results show that our method achieves significant improvements in efficiency, offering superior real-time processing capabilities and reliability compared to existing solutions in ER-on-demand tasks.
1.0.2 Related Work.
@@
Works related to this paper fall under three groupings below.
@@
Graph Constraints. In recent years, graph constraints have emerged as a prominent research area due to their relevance in data quality management. For instance, graph keys (GKeys) [22] leverage graph isomorphism properties to uniquely identify entities within a graph. Graph functional dependencies [23] apply attribute-value dependencies (similar to CFDs [24]) to the topological structures of graphs, while graph entity dependencies (GEDs) [25] further integrate the semantics of both GFDs and GKeys. In this study, we adopt graph differential dependencies (GDDs) [21], an extension of GEDs. GDDs introduce similarity and matching semantics into the existing structure, making them suitable for defining declarative matching rules. This characteristic allows GDDs to better capture approximate matching relationships in graphs, thereby improving the accuracy and robustness of entity resolution (cf. [21, 26, 28]).
@@
Progressive ER. Madhavan et al. [15] were among the first to implement a progressive data integration system in Google Base, demonstrating how web data can be integrated progressively under time and resource constraints. Since then, progressive methods have been widely applied to schema mapping [16, 18, 19, 20] and entity resolution [11, 9], especially in scenarios with limited computational capacity or time for debugging. Existing progressive ER methods employ blocking and scheduling strategies to prioritize high-probability matches for faster discovery [9, 11]. However, these approaches have two key limitations. First, since result emission is not performed at the entity level, matching remains incomplete, potentially leading to incorrect intermediate results that fail to meet user requirements. Second, it lacks support for graph data structures, e.g., entity profiles generated from partial cluster matches may result in inaccurate outcomes.
@@
On-Demand ER. The most related works to this paper are in [14, 10, 17]. Query-Driven Approach (QDA) [14] streamlines data cleaning by analyzing selection predicates within a block to reduce comparisons. However, its lack of progressive execution capability makes it unsuitable for real-time applications. On the other hand, BrewER [10] incrementally cleans data during SQL SP query evaluation using prioritized attribute comparisons. While similar to our approach, it has three key limitations: it requires computationally expensive Cartesian products to generate candidates, demands whole-dataset blocking preparation, and lacks support for graph relationships. QueryER [17] addresses ER in the relational data setting through user-specified queries, operating within a well-defined schema using a dedicated parser, planner, and executor. Its goal is to integrate ER capabilities into query engines. In contrast, our work presents a standalone, on-demand ER solution for property graphs, without requiring integration into query processing systems.
@@ 2 Preliminaries
This section presents relevant concepts and definitions used in the paper.
@@ 2.1 Basic Definitions and Notions
An entity profile for a real-world entity is a tuple $p=\langle pid,eid,type,P,R\rangle$, where pidpid is a unique profile identifier, eideid is the (possibly unknown) real-world entity ID, and typetype denotes its category (e.g., person, location). $P$ is a set of attribute-value pairs $(A,c)$, and $R$ a list of relations $(rela,pid^{\prime})$ linking the entity to other profiles. If two profiles, $p_{1},p_{2}$, represent the same entity, then $p_{1}.eid=p_{2}.eid$.
@@ 2.2 Graph Differential Dependency (GDD)
Consider $\varphi:(Q_{1}[x,x^{\prime},y],\Phi_{X}\to\Phi_{Y})$, where $\Phi_{X}=\{\delta_{Age}(x,x^{\prime})\leq 3\wedge\delta_{Ph}(x,x^{\prime})=0\}$ and $\Phi_{Y}=\{\delta_{eid}(x,x^{\prime})=0\}$. This states that if two users in $Q_{1}$ have an Age difference within 33 and identical Phone numbers, they refer to the same real-world entity. ∎
@@ 4  $edgeWeight\leftarrow\sum_{r\in\mathcal{R}}\text{if }x,x^{\prime}\text{ satisfy }r$
6     Add edge $(x,x^{\prime})$ with edgeWeightedgeWeight to BlockingGraphBlockingGraph
@@ 12        Compare($v,candidate$), updateMatchLists($v,candidate$)
In addition, FastER offers interpretability by design. The filtering process is driven by user-defined GDD rules, which specify explicit graph patterns and attribute constraints. Each candidate match retained by the system can be traced back to the rules it satisfies, allowing users to understand and refine the matching behavior based on domain semantics.
Figure 3: The overall architecture of FastER. (a) The dirty property graph to be cleaned, (b) Subgraph obtained after graph pattern filtering, (c) Subgraph further refined by constraint-based filtering, (d)Constraints in GDD, (e) Blocking graph generated based on candidate matches and constraints, (f) The sorted profile list, (g) The neighborhood of node V11 within the same block, (h) The comparison list after processing V11 during the emission phase, (i) Comparison results of V11’s related candidates.
@@ 3.2 Constraints-Based Filtering
We define several distance constraints, which require that the differences between node pairs on certain attributes must fall within specific thresholds, such as last names, phone numbers, and ages. In this example, we define four constraints: - Lastname (LN) edit distance: $\delta_{\text{LN}}(x,x^{\prime})\leq 3$ - Firstname (FN) edit distance: $\delta_{\text{FN}}(x,x^{\prime})\leq 3$ - Phone (PH) edit distance: $\delta_{\text{PH}}(x,x^{\prime})\leq 3$ - Age (AGE) difference: $\delta_{\text{AGE}}(x,x^{\prime})=0$. Only candidate pairs that satisfy at least one of these constraints will be retained for further processing.
@@
3.3 Blocking Graph Construction and PPS
Next, we construct a Blocking Graph based on the refined candidate list. In this graph, nodes represent entities, and edges represent potential matches based on their attributes. The weight of each edge is determined by the similarity between connected nodes. For simplicity and clarity, we assume that if two nodes satisfy one rule, the edge weight increases by one. We can define a Blocking Graph as shown in Fig. 3 e). This graph naturally divides the nodes from the original subgraph into two blocks, reducing the number of node comparisons and improving matching efficiency.
@@ 3.3 Blocking Graph Construction and PPS
Progressive Profile Scheduling (PPS). Once the Blocking Graph is constructed, we adopt the Progressive Profile Scheduling (PPS) technique to efficiently perform entity matching. PPS is an entity-centric scheduling method designed to prioritize nodes with the highest likelihood of matching by assessing their duplication probability.
@@
3.4 Time Complexity
Let $G=(V,E)$ be a property graph with $|V|=n$ nodes and $|E|=m$ edges, and let $Q=(V_{Q},E_{Q})$ be a small graph pattern. Subgraph isomorphism using Cypher (the query language used by Neo4j) in Neo4j has worst-case time complexity $O(m\cdot|E_{Q}|)$, which reduces to $O(m)$ for constant-size patterns (as discussed in Appendix A.1, Section 1).
Let $C$ be the number of candidate node pairs identified by pattern matching. Each of these undergoes $d$ GDD constraint checks, giving $O(C\cdot d)$ time. Since $d$ is typically constant, this simplifies to $O(C)$. After constraint filtering, only a fraction $\beta\in(0,1)$ of the pairs remain. Let $C^{\prime}=\beta C$ denote the number of valid pairs after filtering. Blocking graph construction on these $C^{\prime}$ pairs requires $O(C^{\prime}\cdot r)$ for edge weight computation, where $r$ is the number of features per edge, and $O(C^{\prime}\log C^{\prime})$ for sorting by similarity score. Finally, Progressive Profile Scheduling (PPS) incrementally compares each query-relevant profile (total $N$) to its $k^{\prime}$ most promising candidates (typically $k^{\prime}\ll C^{\prime}$), yielding $O(N\cdot k^{\prime})$ complexity for the matching phase.
Summing the steps, the total complexity is: $O(m+\beta C+C^{\prime}\log C^{\prime}+N\cdot k^{\prime}).$ As shown in Appendix A.1, $\beta$ and $k^{\prime}$ are empirically small in most cases (e.g., $\beta<0.1$, $k^{\prime}<5$), yielding an overall complexity close to linear in the graph size $m$ and number of target entities $N$.
@@ 4 Experimental Evaluation
This section empirically evaluates our proposal w.r.t. different SOTA solutions.
4.1 Experimental Setting
All experiments were performed on a PC with a 2.40GHz Intel Xeon Silver 4210R processor and 32GB of memory, running Linux OS. The experiments were implemented in Python 3.10 and interacted with the Neo4j graph database (version 5.19.0).
Datasets. In this study, we used five graph benchmark datasets and five relational benchmark datasets (cf. Table 1) from various domains. All datasets used in this paper are publicly available on the GitHub repository [^1] https://github.com/Zaiwen/On_Demand_Entity_Resolution. Although FastER is specifically designed for property graphs, we show it can also be applied in the relational data setting. Thus, we transformed the relational datasets in Table 1 into graph representations by leveraging their attribute relationships. The detailed conversion methods are provided in the  GitHub repository. The reported values for Nodes, Edges, Node Types, and Edge Types in the relational datasets are based on their graph-transformed statistics. We apply automated GDD mining algorithms [21, 27], followed by manual selection of a small number of high-quality rules to ensure interpretability and relevance. This process is detailed in subsection 4.6. To ensure a fair and consistent comparison across methods, all approaches use the same oracle matcher.
Evaluation Metrics. It is important to note that in our experiments, we do not directly evaluate precision or F1-score. FastER is a matcher-agnostic framework that focuses on filtering and scheduling components. Therefore, we adopt an oracle matcher—i.e., entity pairs are labeled using ground truth— to isolate the effects of the matching function and enable a fair comparison of the framework’s efficiency and recall capabilities. Additionally, following standard practice in ER on-demand literature [10], the final matching confirmation stage guarantees a precision of 1. As a result, we exclude precision and F1-score from our evaluation. Instead, we focus on metrics such as query recall and error rate, which are discussed in detail later. We explore the following questions in subsequent subsections:
  1. 1.
    How does FastER perform in ER-on-demand scenario (Exp-1)?
  2. 2.
    How does FastER compare to SOTA batch processing methods (Exp-2)?
  3. 3.
  4. 4.
    How does FastER perform on different components (Exp-4)?
  5. 5.
    How can users be guided to select appropriate rules (Exp-5)?